# Reliable and efficient inverse analysis using physics-informed neural networks with distance functions and adaptive weight tuning

### v1

@@ Abstract
Physics-informed neural networks have attracted significant attention in scientific machine learning for their capability to solve forward and inverse problems governed by partial differential equations. However, the accuracy of PINN solutions is often limited by the treatment of boundary conditions. Conventional penalty-based methods, which incorporate boundary conditions as penalty terms in the loss function, cannot guarantee exact satisfaction of the given boundary conditions and are highly sensitive to the choice of penalty parameters. This paper demonstrates that distance functions, specifically R-functions, can be leveraged to enforce boundary conditions, overcoming these limitations. R-functions provide normalized distance fields, enabling accurate representation of boundary geometries, including non-convex domains, and facilitating various types of boundary conditions. We extend this distance function-based boundary condition imposition method to inverse problems using PINNs and introduce an adaptive weight tuning technique to ensure reliable and efficient inverse analysis. We demonstrate the efficacy of the method through several numerical experiments. Numerical results show that the proposed method solves inverse problems more accurately and efficiently than penalty-based methods, even in the presence of complex non-convex geometries. This approach offers a reliable and efficient framework for inverse analysis using PINNs, with potential applications across a wide range of engineering problems.
Keywords scientific machine learning, physics-informed neural network, implicit distance function, adaptive weight tuning, inverse analysis, incompressible flow
@@
The main contributions of our study can be summarized as follows:
  • First application of R-function-based distance functions to inverse problems using physics-informed neural networks.
  • Synergistic integration of distance functions and adaptive weight tuning for reliable inverse analysis.
  • Enhanced accuracy and efficiency in inverse problem solutions across diverse and complex geometries.
@@ 1 Introduction
In recent years, neural networks have emerged as an alternative approach for solving PDEs, with the pioneering studies appearing in the 1990s [1, 2, 3, 4]. With recent advancements in computational hardware and the spread of machine learning libraries, neural network-based methods have further gained significant attention across diverse disciplines, not only in computer vision and natural language processing, but also in science and engineering. However, a major challenge associated with neural network-based methods is their lack of explainability and trustworthiness [5, 6]. Such concerns have led to a shift towards the development of models that incorporate prior knowledge, thereby improving their interpretability and reliability compared to purely data-driven methods [7]. This is also the case when applying ML to physical problems, which has led to the emergence of a new research field known as scientific machine learning (SciML) [8]. Several approaches have been proposed to integrate known physical laws into the training process of neural networks. Notable examples include the Physics-Informed Neural Network (PINN) [9], in which the governing PDEs are introduced into the loss function; the Deep Ritz Method [10, 11, 12], where the corresponding energy functional is minimized; and the Variational PINN [13, 14], the Weakly Adversarial Network [15], and the deep mixed residual method [16], which are similar to PINN but relax the regularity of trial solutions by using weak formulations or auxiliary variables. One of the key advantages of PINN and its variants is their flexibility in handling both forward and inverse problems in a unified framework with minimal code modifications. In addition, embedded physical laws serve to effectively constrain the search space of parameters, leading to more accurate solutions and robustness to noisy data [17, 18, 19]. PINNs and variants also reduce the need for large amounts of data, making them suitable for problems with limited data availability [20, 21, 22]. In this regard, PINN can be viewed as a class of Sobolev training networks [23, 24], while one can explicitly introduce an additional regularization [25, 26, 27]. This prior knowledge-based regularization has been effectively utilized in various fields, as demonstrated by Kissas et al. [21] for pressure reconstruction from scarce and noisy medical data, and Sun et al. [28] for airfoil design optimization, to name a few. There are also studies to improve the robustness and approximation capabilities of PINN by introducing architecture modifications [29, 30, 31], basis transformations [32, 33, 34], and domain decompositions [35, 36, 37, 38].
@@
Despite the considerable potential and advancements, the accuracy of PINN solutions is still limited by the treatment of boundary conditions. In practice, many approaches only consider the BCs via the loss function, which we refer to as ‘soft imposition’ of boundary conditions. Although this penalty-based method is relatively easy to implement and has been widely adopted, it can lead to poor approximations if not applied carefully [39]. Potential countermeasures have been proposed, including:
@@
    Increasing the weight of the boundary condition loss to enhance soft imposition effect [10, 40, 41]
@@
    Utilizing multiple networks, each learning the prescribed BCs and the governing PDEs separately [42, 43, 44]
@@
    Enforcing boundary conditions using distance functions to the boundary [2, 39, 45, 46]
For methods (1) and (2), however, the accuracy of the solution is heavily dependent on the choice of the weight (penalty parameter) and the optimization method. In particular, for method (1), several studies have proposed adaptive weight tunings based on statistics or norm of backpropagated gradients [24, 30, 47], or Neural Tangent Kernel theory [48, 49]. Nonetheless, including method (2), these approaches only provide approximations, and therefore the resulting solutions are not guaranteed to satisfy the given BCs. Method (3), on the other hand, is independent of the optimizer or approximator, and is the focus of this paper, referred to as ‘hard imposition’. In this method, the approximate solution is derived by projecting the output of the neural network onto the space that exactly satisfies the boundary conditions. This transformation is performed by using distance functions (in a Euclidean sense), which is constructed based on the theory of R-functions [50, 51, 52]. R-functions are a class of functions to encode logical operations into real-valued functions, and have been utilized in geometric modeling [52, 53, 54]. While the application of R-function-based distance functions to numerical and machine learning methods has been explored for forward problems  [55, 56, 57, 46, 58], their use in inverse problems with PINN has not been fully investigated. In this paper, we extend the method discussed in [59, 51, 52, 46, 58] to address inverse problems using PINN, while also applying an adaptive weight tuning [47]. To the best of our knowledge, this is the first study to apply R-function-based distance functions to inverse problems using PINNs. We demonstrate the effectiveness of this approach through numerical experiments, and show that the proposed method can solve inverse problems more accurately and efficiently than the conventional penalty-based method. The remainder of this paper is organized as follows. In Section 2, we provide an overview of the physics-informed neural network (PINN) and describe the properties of the R-function-based distance function. Next, we discuss how to apply the distance function to enforce boundary conditions in PINN, and introduce an adaptive weight tuning along with the technique to ensure the reliability of inverse analysis using PINN. Consequently, in Section 3, we present a series of numerical experiments to evaluate the performance of the proposed method. Finally, we conclude the paper in Section 4, highlighting the significance of our work.
@@ 2 Methods
2.1 Physics-informed neural network
For the sake of clarity, we first provide a brief overview of the physics-informed neural network (PINN). Consider the following boundary value problem of the Poisson equation as a model problem:
@@ 2.1 Physics-informed neural network
Following [9], we proceed by approximating the solution $u(\bm{x})$ with a multi-layer perceptron (MLP) $\hat{u}(\bm{x};\bm{\theta})$ [60, 61], where $\bm{\theta}$ denotes the set of trainable parameters. Given an input $\mathbf{x}\in\mathbb{R}^{f^{(0)}}$ and an output $\hat{\mathbf{y}}\in\mathbb{R}^{f^{(L)}}$, the forward pass $\mathbf{z}^{(l)}\in\mathbb{R}^{f^{(l)}}(l=1,\ldots,L)$ is given by:
@@
where $\mathbf{z}^{(0)}=\mathbf{x}$ and $\mathbf{z}^{(L)}=\hat{\mathbf{y}}$. $\sigma^{(l)}(\cdot)$ is the element-wise nonlinearity, and $\mathbf{W}^{(l)}\in\mathbb{R}^{f^{(l)}\times f^{(l-1)}}$ and $\mathbf{b}^{(l)}\in\mathbb{R}^{f^{(l)}}$ represent the learnable weights and biases, respectively, with $\bm{\theta}={\mathbf{W}^{(l)},\mathbf{b}^{(l)}}_{l=1}^{L}$. In this context, $\mathbf{x}$ corresponds to the spatial coordinates $\bm{x}$, and $\hat{\mathbf{y}}$ represents the approximate solution $\hat{u}$. PINN is trained by minimizing the following loss function:
@@
where each of the terms is evaluated via (quasi-)Monte Carlo approximation (e.g., [62]):
[…]
Here, $\lambda_{\text{BC}},\lambda_{\text{Data}}\in\mathbb{R}_{\geq 0}$ are weights (penalty parameters) to control the contributions of boundary conditions and observed data to the loss function. The terms $\mathcal{L}_{\text{PDE}}$, $\mathcal{L}_{\text{BC}}$, and $\mathcal{L}_{\text{Data}}$ represent the loss functions associated with the PDE residual, BC, and observed data, respectively. For forward problems, $\lambda_{\text{Data}}$ can be set to zero, whereas for inverse problems, it is assigned a positive value to reflect the observed data. The penalty parameter $\lambda_{\text{BC}}$ is often set to a large value to better propagate the effects of boundary conditions [10, 41, 63]. These weights may be predefined as fixed constants or adaptively adjusted during the training process [24, 30, 47, 49, 31].
2.2 R-function-based distance function
Let $\Phi(\bm{x})$ be the exact distance function (EDF) from a point $\bm{x}\in\Omega$ to the boundary $\Gamma=\partial\Omega$, and $\phi(\bm{x})$ represent the approximate distance function (ADF). By definition, $\Phi$ satisfies the following properties [64]:

### v3

# Reliable and efficient inverse analysis using physics-informed neural networks with normalized distance functions and adaptive weight tuning

@@ Abstract
Physics-informed neural networks have attracted significant attention in scientific machine learning for their capability to solve forward and inverse problems governed by partial differential equations. However, the accuracy of PINN solutions is often limited by the treatment of boundary conditions. Conventional penalty-based methods, which incorporate boundary conditions as penalty terms in the loss function, cannot guarantee exact satisfaction of the given boundary conditions and are highly sensitive to the choice of penalty parameters. This paper demonstrates that distance functions, specifically R-functions, can be leveraged to enforce boundary conditions, overcoming these limitations. R-functions provide normalized distance fields, enabling flexible representation of boundary geometries, including non-convex domains, and facilitating various types of boundary conditions. Nevertheless, distance functions alone are insufficient for accurate inverse analysis in PINNs. To address this, we propose an integrated framework that combines the normalized distance field with bias-corrected adaptive weight tuning to improve both accuracy and efficiency. Numerical results show that the proposed method provides more accurate and efficient solutions to various inverse problems than penalty-based approaches, even in the presence of non-convex geometries with complex boundary conditions. This approach offers a reliable and efficient framework for inverse analysis using PINNs, with potential applications across a wide range of engineering problems.
Keywords scientific machine learning, physics-informed neural network, normalized distance function, adaptive weight tuning, inverse analysis, incompressible flow
@@
The main contributions of this study are as follows:
  • Providing a systematic classification of potential hard imposition approaches in PINNs and clarifying their differences in terms of geometric representation and boundary condition treatment.
  • Employing R-function-based normalized distance fields to impose a wide range of boundary conditions in PINNs over convex and non-convex domains.
  • Integrating normalized distance fields with bias-corrected adaptive weight tuning to achieve faster convergence and accurate inverse analysis in PINNs.
@@ 1 Introduction
In recent years, neural networks have emerged as an alternative approach for solving PDEs, with the pioneering studies appearing in the 1990s [Dissanayake1994, Lagaris1998, Leephakpreeda2002, McFall2009]. With recent advancements in computational hardware and the spread of machine learning libraries, neural network-based methods have further gained significant attention across diverse disciplines, not only in computer vision and natural language processing, but also in science and engineering. However, a major challenge associated with neural network-based methods is their lack of explainability and trustworthiness [Arrieta2020, Roscher2020]. Such concerns have led to a shift towards the development of models that incorporate prior knowledge, thereby improving their interpretability and reliability compared to purely data-driven methods [vonRueden2023]. This is also the case when applying ML to physical problems, which has led to the emergence of a new research field known as scientific machine learning (SciML) [Rackauckas2021SciML]. Several approaches have been proposed to integrate known physical laws into the training process of neural networks. Notable examples include the Physics-Informed Neural Network (PINN) [Raissi2019PINN], in which the governing PDEs are introduced into the loss function; the Deep Ritz Method [E2018DRM, Liao2021DNM, Samaniego2020Energy], where the corresponding energy functional is minimized; and the Variational PINN [Kharazmi2019VPINN, Kharazmi2021hpVPINN], the Weakly Adversarial Network [Zang2020WAN], and the deep mixed residual method [Lyu2022MIM], which are similar to PINN but relax the regularity of trial solutions by using weak formulations or auxiliary variables. One of the key advantages of PINN and its variants is their flexibility in handling both forward and inverse problems in a unified framework with minimal code modifications. In addition, embedded physical laws serve to effectively constrain the search space of parameters, leading to more accurate solutions and robustness to noisy data [Deguchi2023Doboku, Wong2022Robustness, Eivazi2024]. PINNs and variants also reduce the need for large amounts of data, making them suitable for problems with limited data availability [Tartakovsky2020SubsurfaceFlow, Kissas2020MRI, Arzani2021]. In this regard, PINN can be viewed as a class of Sobolev training networks [Czarnecki2017Sobolev, Maddu2022InvDirichlet], while one can explicitly introduce an additional regularization [Yu2022GradEnhc, Son2021Sobolev, Vlassis2021Sobolev]. This prior knowledge-based regularization has been effectively utilized in various fields, as demonstrated by Kissas et al. [Kissas2020MRI] for pressure reconstruction from scarce and noisy medical data, and Sun et al. [Sun2023ShapeOptimization] for airfoil design optimization, to name a few. There are also studies to improve the robustness and approximation capabilities of PINN by various means. These include architecture modifications such as employing gate or residual structures [Sirignano2018DGM, Wang2021GradPath, Wang2023ExpertGuide]; basis transformations with Fourier [Tancik2020FourierFeature, Wang2021FourierFeature] or polynomial series [Tang2023PolyPINN]; and domain decompositions, where distinct neural network approximations are applied to individual subdomains, with predetermined non-overlapping [Jagtap2020XPINN], overlapping [Li2023D3M, Moseley2023FBPINN], or trainable overlapping subdomains [Hu2023APINN].
@@
Despite the considerable potential and advancements, the accuracy of PINN solutions is still limited by the treatment of boundary conditions. In practice, many studies only consider the BCs via the loss function, which we refer to as ‘soft imposition’ of boundary conditions. Although this penalty-based method is relatively easy to implement and has been widely adopted, it can lead to poor approximations if not applied carefully [Sun2020Surrogate]. Potential countermeasures have been proposed, including:
@@
    Increasing the weight of the boundary condition loss to enhance soft imposition effect [E2018DRM, Chen2020DGMvsDRM, Wight2021]
@@
    Utilizing multiple networks, each learning the prescribed BCs and the governing PDEs separately [Berg2018, Rao2020arXiv, Sheng2021PFNN]
@@
    Enforcing boundary conditions using distance functions to the boundary [Lagaris1998, Sun2020Surrogate, Lu2021HardDesign, Sukumar2022MixedBC]
For methods (1) and (2), however, the accuracy of the solution is heavily dependent on the choice of the weight (penalty parameter) and the optimization method. In particular, for method (1), several studies have proposed adaptive weight tunings based on statistics or norm of backpropagated gradients [Maddu2022InvDirichlet, Wang2021GradPath, Deguchi2023DynNorm], or Neural Tangent Kernel theory [Jacot2018NTK, Wang2022NTK]. Nonetheless, including method (2), these approaches only provide approximations, and therefore the resulting solutions are not guaranteed to satisfy the given BCs. Method (3), by contrast, is independent of the optimizer or approximator in terms of boundary condition enforcement, and is therefore referred to as ‘hard imposition’. In this method, the approximate solution is obtained by projecting the output of the neural network onto the space that exactly satisfies the boundary conditions. This projection is realized using distance functions, which can be constructed in several ways. Here, we note that the term ‘hard imposition’ has been used across various contexts, resulting in a broad range of interpretations without clear distinction. In particular, studies employing distance functions, whether normalized or not (as discussed below), or even those approximated by another neural network, which should fall under method (2) above, are sometimes called hard imposition. While we do not attempt to standardize the terminology in this study, it remains of great importance to distinguish the different approaches presented in the literature, so that it is clear which strategy is applicable to a given problem setting. Here, we classify method (3), distance function-based strategy, into the following three distinct classes from two perspectives: flexibility in geometric representation and versatility in boundary condition treatment (see Table 1 for a summary):
  • (3-i)
    Naïve approach: represents the domain as the intersection of half-spaces whose zero level sets are located to coincide with boundary components; applicable only to convex geometries and typically restricted to homogeneous Dirichlet conditions (or at most up to an additive constant for a global translation; presented in, e.g., [Lagaris1998, Lu2021HardDesign]).
  • (3-ii)
    Non-normalized approach: represents the domain boundary as the union of distance functions to boundary components, which are based on R-functions, but joined without normalization; applicable to both convex and non-convex geometries yet still limited to homogeneous Dirichlet conditions (or up to an additive constant)
  • (3-iii)
    Normalized approach: represents the domain boundary as the union of distance functions to boundary components, based on R-functions, and joined with normalization to preserve essential properties near the boundary; applicable to both convex and non-convex geometries, and capable of handling various boundary conditions including spatially varying inhomogeneous Dirichlet and mixed boundary conditions (e.g., [Sukumar2022MixedBC]).
Among them, this study focuses on approach (3-iii), which builds distance functions based on the theory of R-functions [Sheiko1982, Rvachev2001, Shapiro2007], a class of functions to encode logical operations into real-valued functions widely utilized in geometric modeling [Shapiro2007, Shapiro1994, Shapiro1999DiffProperties]. This choice is motivated by the advantages of R-functions in geometric representation and boundary condition enforcement. In comparison to (3-iii), approach (3-i) has been frequently adopted in the literature due to its simplicity. However, it exhibits severe limitations in representing complex geometries and imposing various BCs, while these shortcomings are often overlooked and left undiscussed. Approach (3-ii) also employs R-functions but lacks normalization in the joining process. As a result, its BC enforcement capability is limited, while this class of formulation is readily extendable to (3-iii). The distinctions among these approaches, including normalization properties and associated limitations, are detailed with illustrative examples in Section 2.2.1.
While the application of R-function-based normalized distance fields (approach (3-iii)) to numerical simulations [Shapiro2002SAGE, Tsukanov2003MeshfreeFluid, Millan2015] and machine learning methods [Sukumar2022MixedBC, Berrone2023BCEnforcement, Gladstone2025FOPINN] has been explored for forward problems, their use in inverse problems with PINNs has not been fully investigated. This study addresses the aforementioned limitations in accuracy, attributed to the inadequate treatment of boundary conditions, by employing R-function-based normalized distance fields to enforce various BCs over convex and non-convex domains [Sukumar2022MixedBC, Rvachev2001, Rvachev1995, Shapiro2007]. However, we found that distance functions alone are insufficient to achieve accurate inverse analysis in PINNs. To overcome this limitation, we propose an integrated framework that combines the normalized distance field with a bias-corrected adaptive weight tuning [Deguchi2023DynNorm] to balance contributions from multiple loss components. The proposed integrated approach provides two key benefits. First, normalized distance functions enable precise representation of complex geometries and various types of conditions, thereby improving the reliability of PINN solutions. Second, the adaptive weight tuning provides appropriate weights to multiple objectives from the early stages of optimization, with the help of bias-correction, offering faster convergence and higher accuracy. To the best of our knowledge, this is the first study to extend R-function-based normalized distance functions to inverse analysis using PINNs, further enhanced by integration with bias-corrected adaptive weights. We demonstrate the effectiveness of this integrated approach through numerical experiments, showing improved accuracy and efficiency in solving inverse problems with complex geometries and boundary conditions.
The remainder of this paper is organized as follows. In Section 2, we first provide an overview of physics-informed neural networks (PINNs), taking the Poisson equation as a model problem. We then summarize the fundamental properties of exact and approximate distance functions, which are tightly related to the normalization properties of distance fields. Next, we review the construction of distance fields in approaches (3-i), (3-ii), and (3-iii) and illustrate their differences with examples. We also describe how boundary conditions are incorporated into trial functions. Following this, we introduce a gradient norm-bsaed adaptive weight tuning method, which is necessary for optimizing multiple objectives that arise in inverse problems. Furthermore, we show that adaptive weights can be biased towards the initial values under certain conditions and present a correction strategy to mitigate this issue, along with a technique to further enhance the reliability of inverse analysis. In Section 3, we present a series of numerical experiments to evaluate the performance of the proposed integrated method in comparison with the widely adopted penalty-based approach. Finally, Section 4 concludes the paper by highlighting the significance of our work.
Table 1: Taxonomy of distance function-based hard imposition approaches in the context of PINNs. Classified according to geometric representation flexibility and boundary condition enforcement versatility. See Section 2.2.1 for details.
Naïve (Eq. (13)) Non-normalized (Eq. (14)) Normalized (Eq. (12))
Geom. repr. Convex Convex/non-convex Convex/non-convex
Bound. cond. enf. Hmg. Dirichlet Hmg. Dirichlet Hmg./inhmg. Dirichlet/mixed
Refs. [Lagaris1998, Lu2021HardDesign, Li2024HardAdvcDiff] [Sukumar2022MixedBC, Berrone2023BCEnforcement]
@@ 2 Methods
2.1 Background: physics-informed neural networks
For the sake of clarity, we first provide a brief overview of physics-informed neural networks (PINNs), taking the Poisson equation as a model problem. The discussion can be readily extended to other types of equations or to unsteady problems by employing a space-time strategy, which is a common practice in the literature (see, e.g., [Sun2020Surrogate, Kharazmi2021hpVPINN, Yu2022GradEnhc]).
Consider the following boundary value problem of the Poisson equation:
@@ 2.1 Background: physics-informed neural networks
Following [Raissi2019PINN], we proceed by approximating the solution $u(\bm{x})$ with a multi-layer perceptron (MLP) $\hat{u}(\bm{x};\bm{\theta})$ [Hornik1989, Leshno1993], where $\bm{\theta}$ denotes the set of trainable parameters. Given an input $\mathbf{x}\in\mathbb{R}^{q^{(0)}}$ and an output $\hat{\mathbf{y}}\in\mathbb{R}^{q^{(L)}}$, the forward pass $\mathbf{z}^{(l)}\in\mathbb{R}^{q^{(l)}}(l=1,\ldots,L)$ is given by:
@@
where $\mathbf{z}^{(0)}=\mathbf{x}$ and $\mathbf{z}^{(L)}=\hat{\mathbf{y}}$. $\sigma^{(l)}(\cdot)$ is the element-wise nonlinearity, and $\mathbf{W}^{(l)}\in\mathbb{R}^{q^{(l)}\times q^{(l-1)}}$ and $\mathbf{b}^{(l)}\in\mathbb{R}^{q^{(l)}}$ represent the learnable weights and biases, respectively, with $\bm{\theta}=\{\mathbf{W}^{(l)},\mathbf{b}^{(l)}\}_{l=1}^{L}$. In this context, $\mathbf{x}$ corresponds to the spatial coordinates $\bm{x}$, and $\hat{\mathbf{y}}$ represents the approximate solution $\hat{u}$. PINN is trained by minimizing the following loss function:
@@
where each of the terms is evaluated via (quasi-)Monte Carlo approximation (e.g., [Matsubara2023GLT]):
$\begin{split}\mathcal{L}_{\text{PDE}}&=\int_{\Omega}\left|-\nabla^{2}\hat{u}(\bm{x};\bm{\theta})-f(\bm{x})\right|^{2}\,d\bm{x}\\ &\simeq\frac{1}{N_{\text{PDE}}}\sum_{i=1}^{N_{\text{PDE}}}\left|-\nabla^{2}\hat{u}(\bm{x}_{i};\bm{\theta})-f(\bm{x}_{i})\right|^{2},\end{split}$ (6)
$\begin{split}\mathcal{L}_{\text{BC}}&=\int_{\Gamma_{D}}\left|\hat{u}(\bm{x};\bm{\theta})-g_{D}(\bm{x})\right|^{2}\,d\bm{x}+\int_{\Gamma_{N}}\left|\bm{n}\cdot\nabla\hat{u}(\bm{x};\bm{\theta})-g_{N}(\bm{x})\right|^{2}\,d\bm{x}\\ &\simeq\frac{1}{N_{\text{DBC}}}\sum_{i=1}^{N_{\text{DBC}}}\left|\hat{u}(\bm{x}_{i};\bm{\theta})-g_{D}(\bm{x}_{i})\right|^{2}+\frac{1}{N_{\text{NBC}}}\sum_{i=1}^{N_{\text{NBC}}}\left|\bm{n}\cdot\nabla\hat{u}(\bm{x}_{i};\bm{\theta})-g_{N}(\bm{x}_{i})\right|^{2},\end{split}$ (7)
$\begin{split}\mathcal{L}_{\text{Data}}&=\int_{\Omega}\left|\hat{u}(\bm{x};\bm{\theta})-u_{\text{Data}}(\bm{x})\right|^{2}\,d\bm{x}\\ &\simeq\frac{1}{N_{\text{Data}}}\sum_{i=1}^{N_{\text{Data}}}\left|\hat{u}(\bm{x}_{i};\bm{\theta})-u_{\text{Data}}(\bm{x}_{i})\right|^{2}.\end{split}$ (8)
Here, $\lambda_{\text{BC}},\lambda_{\text{Data}}\in\mathbb{R}_{\geq 0}$ are weights (penalty parameters) to control the contributions of boundary conditions and observed data to the loss function. The terms $\mathcal{L}_{\text{PDE}}$, $\mathcal{L}_{\text{BC}}$, and $\mathcal{L}_{\text{Data}}$ represent the loss functions associated with the PDE residual, BC, and observed data, respectively. For forward problems, $\lambda_{\text{Data}}$ can be set to zero, whereas for inverse problems, it is assigned a positive value to reflect the observed data. The penalty parameter $\lambda_{\text{BC}}$ is often set to a large value to better propagate the effects of boundary conditions [E2018DRM, Wight2021, Rohrhofer2023IEEE]. These weights may be predefined as fixed constants or adaptively adjusted during the training process. The latter includes methods based on gradient statistics [Maddu2022InvDirichlet, Wang2021GradPath], norms [Deguchi2023DynNorm, Wang2023ExpertGuide], or the trace of Neural Tangent Kernel matrices [Wang2022NTK].
2.2 Proposed method: integration of normalized distance and adaptive weight
2.2.1 R-function-based distance fields
Fundamental properties of distance functions
Let $\Phi(\bm{x})$ be the exact distance function (EDF) from a point $\bm{x}\in\Omega$ to the boundary $\Gamma=\partial\Omega$, and $\phi(\bm{x})$ represent the approximate distance function (ADF). By definition, $\Phi$ satisfies the following properties [Biswas2004]: